# Embracing Radiance Field Rendering in 6G: Over-the-Air Training and Inference with 3D Contents

### v1

@@ IV-B2 Algorithmic Model Acceleration
The acceleration of rendering from an algorithmic perspective is also useful to reduce rendering latency. This is especially crucial for NeRF, where synthesizing new views based on a well-trained model requires frequent queries to MLPs for generating density and color at each sampling point. There are generally two approaches from the computer vision community [29, 30, 31]. One approach is to pre-compute and store the neural networks [29], transforming the original NeRF model into more easily accessible data structures to significantly reduce the rendering time. The other approach is to learn the scene features by additionally utilizing other representations like voxel grids [31], allowing for reducing the size of neural networks and leading to less rendering time.
@@ IV-C2 Semantic Communications Enabled Radiance Field Rendering
Under the local-remote co-rendering architecture, both the server and the device collaboratively execute rendering tasks, where frequent data exchanges may incur significant communication overhead. To tackle this issue, semantic communications, a novel communication system design paradigm beyond the conventional bit-level transmission, is becoming an interesting new solution that can be applied to enhance the end-to-end rendering performance. The effective transmission and rendering of 3D contents with semantic communications rely on the semantic knowledge base construction and semantic feature extraction. Specifically, semantic knowledge base provides rich knowledge for semantic information processing (i.e., feature extraction and data recovery) [54], which can be exploited as the side information at both the transmitter and the receiver to reduce the transmitted bits while preserving the QoE. In particular, radiance field models can be exploited as high-quality semantic knowledge base storing fruitful radiance information of 3D scenes [1]. With such knowledge base, the transmitter only needs to extract and transmit essential semantic features to the receiver for real-time 3D reconstruction. Furthermore, to accommodate the dynamic channel conditions especially in the wireless environment, channel condition aware semantic feature selection and transmission policies should be further designed. The semantic communications-based co-inference is appealing for not only reconstructing 3D contents, but also for further performing downstream intelligent tasks such as 3D video analysis. A case study on using semantic communications for 3D human face transmission is presented in the next section to show the value of semantic communications in this direction.
@@ V Case Study: Semantic Communication for 3D Content Transmission
The semantic communications framework with NeRF (NeRF-SeCom) for 3D human face transmission is shown in Fig. 5, which includes face semantic feature extraction, feature selection, feature chunk packing, learning-based feature prediction, and face rendering. We exploit the NeRFBlendShape [58] face NeRF model as the facial knowledge base to assist face semantic feature extraction and face rendering. The NeRFBlendShape face model is expressed as
@@
where $V(\cdot)$ represents the volume rendering. In the following, we provide the detailed workflow of our proposed NeRF-SeCom framework.

### v2

@@ IV-B2 Algorithmic Model Acceleration
The acceleration of rendering from an algorithmic perspective is also useful to reduce rendering latency. This is especially crucial for NeRF, where synthesizing new views based on a well-trained model requires frequent queries to MLPs for generating density and color at each sampling point. This may incur significant time costs in the rendering stage, and thus influence the overall latency of inference. There are generally two approaches from the computer vision community [32, 33, 34]. One approach is to pre-compute and store the neural networks [32], transforming the original NeRF model into more easily accessible data structures to significantly reduce the rendering time. The other approach is to learn the scene features by additionally utilizing other representations like voxel grids [34], allowing for reducing the size of neural networks and leading to less rendering time.
@@ IV-C2 Semantic Communications Enabled Radiance Field Rendering
Under the local-remote co-rendering architecture, both the server and the device collaboratively execute rendering tasks, where frequent data exchanges may incur significant communication overhead. To tackle this issue, semantic communications, a novel communication system design paradigm beyond the conventional bit-level transmission, is becoming an interesting new solution that can be applied to enhance the end-to-end rendering performance. The effective transmission and rendering of 3D contents with semantic communications rely on the semantic knowledge base construction and semantic feature extraction. Specifically, semantic knowledge base provides rich knowledge for semantic information processing (i.e., feature extraction and data recovery) [65], which can be exploited as the side information at both the transmitter and the receiver to reduce the transmitted bits while preserving the QoE. In particular, radiance field models can be exploited as high-quality semantic knowledge base storing fruitful radiance information of 3D scenes [1]. With such knowledge base, the transmitter only needs to extract and transmit essential semantic features to the receiver for real-time 3D reconstruction. Furthermore, to accommodate the dynamic channel conditions especially in the wireless environment, channel condition-aware semantic feature selection and transmission policies should be further designed. The semantic communications-based co-inference is appealing for not only reconstructing 3D contents, but also for further performing downstream intelligent tasks such as 3D video analysis. A case study on using semantic communications for 3D human face transmission is presented in the next section to show the value of semantic communications in this direction.
@@ V Case Study: Semantic Communication for 3D Content Transmission
The semantic communications framework with NeRF (NeRF-SeCom) for 3D human face transmission is shown in Fig. 5, which includes face semantic feature extraction, feature selection, feature chunk packing, learning-based feature prediction, and face rendering. Semantic knowledge base aims to provide rich knowledge for semantic information processing (i.e., feature extraction and data recovery), which is essential in semantic communication systems [65]. We exploit the NeRFBlendShape [69] face NeRF model, which employs a multi-level hash table [33] to associate expression coefficients for accelerating the rendering process. This model serves as the semantic knowledge base, assisting in both face semantic feature extraction and face rendering. The NeRFBlendShape face model is expressed as
@@
where $V(\cdot)$ represents the volume rendering. In our implementation, we first pre-train the NeRFBlendShape-based facial knowledge base with a segment of monocular video of human faces as the training dataset in an offline manner. Next, the pre-trained semantic knowledge base is shared at the transmitter and receiver before 3D face transmission. The semantic knowledge base corresponds to the facial identity, [^3] The research on generalizable face models applicable to different identities is indeed an interesting direction and holds significant potential. How to generate such models is beyond the scope of the current work and can be considered for future research. and as a result only needs to be trained and transmitted once for the same person over a long time period consisting of a large number of video chunks. It is worth noting that the NeRFBlendShape model is lightweight and thus it can be transmitted without influencing the communication latency in real time. In the following, we provide the detailed workflow of our proposed NeRF-SeCom framework.
@@
It is worth emphasizing that although this case study focuses on 3D human face transmission by exploiting the NeRF face model as the semantic knowledge base, our proposed semantic framework can be well generalized to other 3D contents. First, this framework can be implemented in the 3D human avatar transmission by exploiting avatar radiance field models such as [71] as the semantic knowledge base to facilitate semantic information processing. The effectiveness of this paradigm has been demonstrated by the prior works that exploit conventional 3D representations as the semantic knowledge base in human avatar transmission. For example, the work [68] proposed a point cloud semantic communication framework, in which an avatar model storing skeleton and appearance information is exploited as the knowledge base to help semantic information extraction and avatar pose recovery, but this work did not consider radiance field model. Besides, our framework can be implemented in the transmission of complex scenes. We can exploit a radiance field model to store the features of scenes like [50] as the semantic knowledge base, which is a generic model pre-trained on large scene datasets. By leveraging this model at both transmitter and receiver, we only need to transmit features extracted from the images of target scenes at the transmitter and fine-tune the pre-trained scene model based on the received features at the receiver. Accordingly, we can reconstruct the scene following the volume rendering. This is an interesting direction worth investigating in future work.